find delta H reaction if the decomposition of 29.0g of NaCl consumed 125kJ of heat.

2NaCl -> 2Na + Cl2
1
Expert's answer
2020-04-08T13:29:22-0400

n(NaCl) = m/M = 29g/58.5g/mol = 0.5 mol

According to the equation 2 mol of NaCl reacts so:

0.5 mol of NaCl - 125 kJ

2 mol of NaCl - delta H

delta H reaction = 2mol*125kJ/0.5mol = 500 kJ